the mass of Cu2Se formed when 40.0g copper metal reacts with 28.0g selenium



1
Expert's answer
2022-03-07T03:23:02-0500
  • Step 1: Write a balanced equation
  • 2Cu + Se ----> Cu2Se
  • Step 2: Find the moles of copper and selenium in their masses.
  • 64g of copper is in 1 mole, what about 40g
  • 40/64 = 0.625moles of copper.
  • 79g of selenium is in 1 mole, what about 28g
  • 28/70 = 0.4 moles of selenium.
  • Finally, find the excess reactant by using mole ratio.
  • Mole ratio copper: selenium is 2:1
  • 0.625/2=0.3125 moles of copper vs 0.4 moles of selenium.
  • You realize that selenium is in excess.
  • Therefore, we will use the limiting reactant which is copper and mole ratio to get the mass of Cu2Se
  • 0.625moles are rati0 2 what about ratio 1
  • 0.625/2 = 0.3125 moles.
  • You can convert this to mass by multiplying with molar mass of Cu2Se
  • 0.3125 x 206 = 64.375g of Cu2Se